### v1

# VTechAGP: An Academic-to-General-Audience Text Paraphrase Dataset and Benchmark Models

Abstract
Existing text simplification or paraphrase datasets mainly focus on sentence-level text generation in a general domain. These datasets are typically developed without using domain knowledge. In this paper, we release a novel dataset, VTechAGP, which is the first academic-to-general-audience text paraphrase dataset consisting of 4,938 document-level these and dissertation academic and general-audience abstract pairs from 8 colleges authored over 25 years. We also propose a novel dynamic soft prompt generative language model, DSPT5. For training, we leverage a contrastive-generative loss function to learn the keyword vectors in the dynamic prompt. For inference, we adopt a crowd-sampling decoding strategy at both semantic and structural levels to further select the best output candidate. We evaluate DSPT5 and various state-of-the-art large language models (LLMs) from multiple perspectives. Results demonstrate that the SOTA LLMs does not provide satisfactory outcomes, while the lightweight DSPT5 can achieve competitive results. To the best of our knowledge, we are the first to build a benchmark dataset and solutions for academic-to-general-audience text paraphrase dataset.
VTechAGP: An Academic-to-General-Audience Text Paraphrase Dataset and Benchmark Models
1 Introduction
Text generation aims to produce understandable text in human language from various sources of input data. Among them, text-to-text generation remains an important and challenging task with extensive applications such as language translation Ranathunga et al. (2023); Dabre et al. (2020), paraphrase generation Singh and Josan (2022), text simplification Martin et al. (2023), etc.
Existing text simplification and paraphrase generation datasets (shown in Table 5 in the Appendix) mainly focus on sentence-level translation. The recent paragraph-level dataset WikiAuto Jiang et al. (2020), derived from WikiLarge Zhang and Lapata (2017) and Newsela Xu et al. (2015), still suffers from a lack of domain diversity and specification, which are sourced from Wikipedia and news articles. Besides, the objective of these studies on the aforementioned datasets is to simplify the text for children at lower grade levels. Furthermore, existing works on lay summarization involves brief summarization limited to medicine and biological domains Devaraj et al. (2021); Flores et al. (2023); Jiang et al. (2024); Zaman et al. . However, our proposed VTechAGP involves translation from technical language to general-audience language across a broader set of multiple domains. We aim to generate text from the domain-specific to a general level of understanding, while keeping it scientifically accurate and easy to understand, in order to encourage and facilitate interdisciplinary collaboration across different research fields.
In this paper, we pioneer the research of academic-to-general-audience text generation by introducing a new benchmark dataset VTechAGP, which is derived from electronic theses and dissertations (ETDs) at Virginia Tech over twenty-five years. VTechAGP consists of 4,938 document-level abstract pairs (academic abstract and general-audience abstract). VTechAGP also provides other information such as title, discipline, degree level, etc. This auxiliary information shows the potential of VTechAGP for other tasks such as topic generation, etc. In addition, the abstracts in VTechAGP come from multiple domains (colleges) and VTechAGP is labeled with each specific domain and provides the domain knowledge keywords. More details about VTechAGP are presented in Sec. 3 and Table 6.
Based on VTechAGP, we evaluate several SOTA pre-trained large language models (LLMs), such as LLaMA2 Touvron et al. (2023), Claude2 Anthropic (2023), ChatGPT Brown et al. (2020), etc., to establish the baseline performance. However, these SOTA pre-trained LLMs have demonstrated the following limitations: (1) Some LLMs do not provide public APIs, or the APIs are not free. Also, some LLMs (e.g., Claude2) do not provide fine-tuning, making them less adaptable to specific tasks. (2) The model size of LLMs is very large. For example, LLaMA2 has about 65 billion parameters. Fine-tuning these LLMs is resource-intensive in terms of memory and computation time. Even the inference implementation requires more memory and time. (3) The pre-trained LLMs do not show competitive performance for the academic-to-general-audience text paraphrasing task on VTechAGP in Sec. 5.
To address the above challenges, we propose DSPT5, a dynamic soft prompt-based generative model with the crowd sampling decoding strategy during the inference stage. DSPT5 is built based on the pre-trained T5 Raffel et al. (2020), which has only about 220 million parameters. In particular, the dynamic soft prompt template in DSPT5 can automatically adapt to different academic domains by changing the keywords extracted from the academic abstract. The prompt encoder in DSPT5 is trained to generate and fine-tune keyword vectors combined with the dynamic prompt template. To this end, we design a hybrid loss function with generative language model loss and contrastive loss to jointly learn the generated text representations as well as the ability to distinguish technical keywords from non-technical keywords. During inference, DSPT5 employs two alignment functions at both the semantic and structural levels to select the best candidate for the final generated output.
The contributions can be summarized as follows: (1) Dataset: We construct VTechAGP, the first academic-to-general-audience text paraphrase dataset. VTechAGP is a document-level text generation dataset with multiple technical domains. (2) Baselines: We implement several SOTA LLMs as benchmarks to compare the performance with our proposed model DSPT5. Experimental results show that there is still a huge room for further improvement of the existing LLMs. (3) Approach: We propose a lightweight model, DSPT5, which utilizes dynamic soft prompts with a hybrid loss function and a new crowd decoding strategy. Experimental results show that DSPT5 can achieve competitive results with SOTA LLMs. (4) Evaluation: We explore various evaluation metrics for the academic-to-general-audience text paraphrasing task on VTechAGP from different perspectives, including document-level embedding-based, word-based, and end-to-end metrics. In addition, simplicity, diversity, readability, and toxicity are also considered for the performance evaluation.
2 Related Work
2.1 Text Generation Datasets
The most commonly used datasets for text generation focus on cross-lingual machine translation, such as WMT datasets Farhad et al. (2021); Bojar et al. (2016) for news translation, IWSLT datasets Scarton et al. (11 2-3 2019); Lee et al. (2022) for document translation of TED talks, parallel corpus datasets (Europarl Koehn et al. (2003), UN Parallel Corpus Ziemski et al. (2016), OPUS Zhang et al. (2020), and Tatoeba Tiedemann (2020)) for sentence-level translation. Additional text generation datasets focus on text/sentence simplification, such as datasets from news articles Xu et al. (2015); Stodden et al. (2023), clinical reports Luo et al. (2022a), Wikipedia Zhu et al. (2010); Xu et al. (2016a); Zhang and Lapata (2017); Alva-Manchego et al. (2020); Naderi et al. (2019); Aumiller and Gertz (2022), and other language learning resources Vajjala and Lučić (2018a). Existing paraphrase datasets include phrasal and lexical paraphrases Ganitkevitch et al. (2013), question pairs from Wikidata Fader et al. (2013), and Quora Wang et al. (2017), the same posts shared by Twitter Lan et al. (2017), different image captions Lin et al. (2014), and a machine translation dataset involving back translation Huang et al. (2023). However, these datasets are limited to sentence-level translation and non-academic data with a lack of domain expert knowledge.
2.2 Text Generation Methods
Table 1: Dataset statistics of VTechAGP over eight colleges. Statistics are reported in the format of academic (source) / general audience (target) documents.
College # Doc. Avg. # Sentence Avg. Sentence Len. MTLD Readability Consensus
Agriculture and Life Sciences (ALS) 429 15.06 / 13.80 27.42 / 26.37 73.59 / 68.58 15th-16th / 14th-15th
Architecture, Arts, and Design (AAD) 226 10.83 / 8.99 26.67 / 26.07 68.24 / 68.76 13th-14th / 10th-11th
Engineering (ENG.) 2451 14.03 / 11.99 26.87 / 25.66 69.93 / 66.47 13th-14th / 14th-15th
Liberal Arts and Human Sciences (LAHS) 675 10.12 / 8.60 30.24 / 28.46 68.02 / 67.06 16th-17th / 15th-16th
Natural Resources and Environment (NRE) 269 14.48 / 12.66 28.87 / 27.33 76.69 / 71.99 15th-16th / 10th-11th
Science (SCI.) 679 14.19 / 11.75 27.35 / 25.73 74.90 / 68.48 16th-17th / 15th-16th
Business (BUS.) 63 11.76 / 10.24 28.55 / 27.63 68.46 / 70.35 18th-19th / 17th-18th
Veterinary Medicine (VM) 146 15.61 / 13.16 26.83 / 25.92 77.77 / 70.15 16th-17th / 14th-15th
All 4938 13.68 / 11.66 27.68 / 26.38 71.36 / 67.91 15th-16th / 13th-14th
With the emergence of Transformers Vaswani et al. (2017), pre-trained language models (PLMs) – such as BERT Devlin et al. (2019), BART Lewis et al. (2020), ChatGPT Brown et al. (2020), T5 Raffel et al. (2020), Claude2 Anthropic (2023), and LLaMA2 Touvron et al. (2023) – have been developed and fine-tuned for various generation tasks. For example, fine-tuning LLMs on a customized dataset Sun et al. (2023); Yermakov et al. (2021); Dathathri et al. (2020) with prompts Luo et al. (2023); Wan et al. (2023); Yang et al. (2023); Kew et al. (2023); Chen et al. (2023) is a popular method for text generation. Several decoding strategies (i.e, minimum Bayes decoding Suzgun et al. (2023), beam search Yoon and Bak (2023); Zhang et al. (2020), probability-based sampling Li et al. (2022); Guo et al. (2018); Xu et al. (2022), nucleus sampling Holtzman et al. (2020), prefix-adaptive decoding Pei et al. (2023), and contrastive loss An et al. (2022)) are applied to select the best generated candidates. However, these studies only focus on paraphrase generation or text simplification in a general domain. Our focus is not on shortening the text, but rather on how these PLMs can paraphrase academic language into non-academic language while maintaining accuracy and simplicity.
3 Dataset Construction
3.1 Data Source and Dataset Collection
Virginia Tech has been a leader in electronic theses and dissertations (ETDs) for more than twenty-five years. It was the first university to require electronic submission of ETDs, beginning in 1997. Virginia Tech’s ETDs are accessible through VTechWorks, [^1] https://vtechworks.lib.vt.edu/ a digital repository created through a collaboration between the Graduate School and the University Libraries. In the fall of 2016, the Graduate School added a new requirement for ETDs: the inclusion of a general audience abstract in addition to the traditional academic abstract. The ETD submission system was updated in 2019 to include a separate field for the general audience abstract. Since then, most ETDs include both an academic abstract and a general audience abstract, which are captured as distinct metadata fields in VTechWorks. Like many institutional repositories, VTechWorks supports the Open Archives Initiative Protocol for Metadata Harvesting (OAI-PMH) Lagoze et al. (2002). OAI-PMH is a standard protocol for retrieving metadata records from digital repositories. It provides a framework that enables data exchange between various systems, supporting consistent metadata formats. To create the VTechAGP dataset, we used OAI-PMH to harvest metadata from VTechWorks. We specifically queried the VTechWorks OAI-PMH endpoint to retrieve metadata records for ETDs containing both academic and general audience abstracts. The OAI-PMH endpoint provided us with an XML record for each ETD. This data encompassed a range of metadata elements defined in a qualified Dublin Core schema. We identified specific fields necessary for our dataset and extracted the text content. This step involved parsing the XML structure, locating the relevant elements, and retrieving the textual data they contained. We mapped the extracted metadata to specific columns in a CSV file. Each piece of extracted metadata was mapped to a corresponding column header in the CSV file: ‘identifier_uri’, ‘title’, ‘abstract’, ‘abstract_general’, ‘subject_terms’, ‘discipline’, ‘department’, ‘degree’, ‘degree_level’, and ‘type’. A description is given in Table 6 in the Appendix. We provide the data analysis in Sec. B.
3.2 Task Definition
Figure 1: Dynamic soft prompt generation of our proposed model DSPT5. The framework includes a prompt encoder and a generative language model fine-tuning together with a hybrid loss.
We name the dataset VTechAGP and propose the academic to the general audience text paraphrase task. Let $D=\left\{X_{n},Y_{n},A_{n}\right\}_{n=1}^{N}$ denote a dataset where $X_{n}$ is the source document, $Y_{n}$ is the target document, $A_{n}$ is the auxiliary information (subject terms, which are the keywords in theses/dissertations provided by the author) of the source document, and $N$ is the number of documents. Given an input sequence of words $X=[x_{1},\cdots,x_{S}]$ with length $S$ and auxiliary information $A=[a_{1},\cdots,a_{A}]$ with length $A$, we aim to generate an output sequence of words $Y=[y_{1},\cdots,y_{T}]$ with length $T$ that retains the original meaning as $X$, but reduces the complexity to improve readability and comprehension without domain knowledge. In this task, our goal is to find the set of model parameters that maximize $\prod_{n=1}^{N}p_{model}(Y_{n}|X_{n})$.
4 Approach
Figure 1 shows our proposed DSPT5 with two main components: dynamic soft prompt generation with hybrid loss and crowd sampling decoding strategy.
4.1 Backbone
A generative language model prescribes the generation of a sentence as a sequence of word predictions based on context. T5 Raffel et al. (2020) has shown high performance and few-shot abilities on various language understanding tasks, including text simplification and language translation Brown et al. (2020). The decoding capacity of the generative language model can help to generate arbitrary content well when given appropriate prompts. Therefore, we use the pre-trained version of T5 to initialize our model DSPT5. We fine-tune T5 with dynamic prompts on the training set of VTechAGP  [^2] FLAN-T5 shows a worse performance based on our task. During the inference stage, we implement the crowd sampling decoding strategy to better select word candidates for text generation.
4.2 Dynamic Soft Prompt Generation
In our proposed DSPT5, we introduce an automatic customized soft prompt generation process that includes three main steps: (1) Dynamic Prompt Generation, (2) Soft Prompt Encoder, and (3) Model Training with Hybrid Loss.
4.2.1 Dynamic Prompt Generation
To better control the output generated by T5 for our task, we added a prompt template before the source input (academic abstracts). Instead of using a static prompt template, such as "Generate another version of the provided document for general audiences", we design a dynamic prompt template that can learn to generate academic keywords to adapt to different academic domains. As opposed to adding the keywords directly in front of the prompt Blinova et al. (2023), we designed a template (which is used in the experiments): "Generate the document by reducing the domain knowledge of " + keywords, where the keywords (auxiliary information) $K^{pos}=[k^{pos}_{1},\cdots,k^{pos}_{N}]$ are generated by KeyBERT Grootendorst (2020) 333 https://github.com/MaartenGr/KeyBERT and sorted by their importance score:
${}K^{pos}=sorted(KeyBERT(X_{i}))$ (1)
$N$ is the number of keywords extracted from the source text by KeyBERT. Since KeyBERT does not support fine-tuning on custom datasets, we develop a soft prompt encoder in Sec. 4.2.2 to further fine-tune the embeddings of $K^{pos}$ by minimizing the distance between the embeddings of $K^{pos}$ and the embeddings of the subject terms from the auxiliary information $A$ (golden label). Therefore, dynamic prompts can be automatically generated using the keywords extracted from KeyBERT, and the fine-tuned embeddings of such keywords can be obtained by the soft prompt encoder in the inference stage. Details of loss are introduced in Sec. 4.2.3.
4.2.2 Soft Prompt Encoder
We propose a prompt encoder aiming to generate a soft prompt representing the hidden representation of the text to be generated. The prompt generator can be any sequence-to-sequence model. We use the T5 Raffel et al. (2020) encoder as the backbone for the prompt encoder. We first obtain the embeddings $E_{prompt}=[e_{1},\cdots,e_{A}]$ and $E_{X}=[e_{1},\cdots,e_{S}]$ of a given sequence of tokens $P=[p_{1},\cdots,p_{A}]$ and $X=[x_{1},\cdots,x_{S}]$ from the partial fixed prompt and the source text, respectively. Next, we get the output representations of the prompt encoder $r_{i}$, which can be formulated as:
[…]
where $f_{\phi}$ is the prompt (T5) encoder, $h_{i}$ is the last hidden state of the T5 encoder, and $A_{i}$ has the keywords of the subject terms. A fully connected layer is added after the T5 encoder, where $W\in\mathbb{R}^{n\times d}$ and $b\in\mathbb{R}^{n}$ are trainable. $n$ and $d$ are the length and dimension of the keywords. Similarly, we get the representations of the extracted keywords $r^{pos}_{i}$:
[…]
where $K^{pos}$ is from Equ. 1. We construct negative samples to cover non-academic words. As shown in Figure 1, we implement a sampler that randomly samples words from the source text, where the embeddings of the negative samples $r^{neg}_{i}$:
[…]
where $X_{i}$ is the input source document. The checker in Figure 1 ensures that $K^{neg}\cap K^{pos}=\varnothing$.
4.2.3 Model Training with Hybrid Loss
Given the representations of the extracted soft keyword prompts $r^{pos}_{i}$, the embeddings of the prompt template $E_{prompt}$, and the source input $E_{X}$, the output target text generated by our proposed model:
[…]
where $GLM$ is the T5 backbone, $\theta$ is the model parameter, $\oplus$ denotes concatenation, and $\hat{y}$ is the generated general audience document. DSPT5 is trained to fit the mapping from academic abstracts to general audience abstracts. Formally, let the dataset be $D$ with size $N$. DSPT5 aims to maximize the standard log-likelihood of the target documents over all training samples of $D$:
${}\mathcal{L}_{ce}=\sum_{n}^{N}\sum_{s}^{S_{n}}logP_{\theta}(y_{n,s}|y_{n,<s},x_{n})$ (6)
where $y_{n,s}$ is the s-th word of the general audience abstract in the n-th sample, $S_{n}$ is the length of the target output $y_{n}$, and $\theta$ gives the parameters of DSPT5. To further improve the performance of DSPT5, we add an additional contrastive loss that forces the prompt encoder to generate the embeddings of extracted keywords from KeyBERT $r^{pos}$ that can be more similar to the representations of subject terms $r^{key}$ provided from the source data, while steering away from negative extracted words $r^{neg}$, to encourage the model to bring $r^{key}$ and $r^{neg}$ closer in the learned feature space, while pushing dissimilar instances $r^{key}$ and $r^{neg}$ apart. Therefore, we propose a new hybrid loss function consisting of the cross-entropy loss in Equ. 6 and a contrastive loss infoNCE van den Oord et al. (2018) $\mathcal{L}_{nce}$:
$\mathcal{L}_{nce}=-log\frac{exp(\frac{r^{key}\cdot r^{pos}}{\tau})}{exp(\frac{r^{key}\cdot r^{pos}}{\tau})+\sum_{r^{neg}}^{N^{neg}}exp(\frac{r^{key}\cdot r^{neg}}{\tau})}$ (7)
where $r^{key}$ is from Equ. 2, $r^{pos}$ is from Equ. 3, $r^{neg}$ is from Equ. 4, $N^{neg}$ is the number of negative samples, and $\tau$ is a temperature hyperparameter. Details in Appendix Sec. G. The final hybrid loss function for DSPT5 model combines $\mathcal{L}_{ce}$ and $\mathcal{L}_{nce}$:
${}\mathcal{L}_{hybrid}=(1-\lambda)\mathcal{L}_{ce}+\lambda\mathcal{L}_{nce}$ (8)
where $\lambda$ is a hyperparameter for $\mathcal{L}_{nce}$.
4.3 Crowd Sampling Decoding
Figure 2: The decoding strategy of crowd sampling.
The decoding objective during inference is to select the best candidate $\hat{y}$ among all possible output candidates. We first implement a stochastic sampling method (temperature sampling Ackley et al. (1985)), which selects the next token by sampling from the truncated distribution to generate multiple output candidates. Expanding on Equ. 5:
$\tilde{P_{\theta}}(\cdot|x)=\frac{exp(P_{\theta}(\cdot|x)/\tau)}{\sum exp(P_{\theta}(\cdot|x)/\tau)}$ (9)
where $\tau\in(0,1]$ is the temperature parameter. Crowd sampling shows significant performance improvement over standard sampling methods across a wide range of open-ended text generation tasks Suzgun et al. (2023). To select the final generated output $\hat{y}$, we implement crowd sampling as shown in Figure 2 to get the candidate that can maximize the sum of alignments with the whole crowd by comparing each candidate to the other candidates using the alignment functions. Given a collection of candidate documents $\mathcal{C}$, we get:
[…]
where $\gamma$ is a hyperparameter to adjust the weight for different alignment functions. To measure both the semantic and structural equivalence of the two texts, we consider BERTScore Zhang* et al. (2020) and Levenshtein similarity Levenshtein (1965) as the alignment functions. BERTScore computes the cosine similarity between candidates, considering their representations as vectors. Levenshtein similarity measures the differences between sequences of tokens in the candidates.
5 Experiments
5.1 Experimental Setup
We evaluate the proposed model DSPT5 over eight different colleges in VTechAGP. The dataset is randomly divided into training and testing sets with a ratio of 0.8:0.2 for all other colleges. We discuss the detailed hyperparameters and configurations of DSPT5 in Appendix Sec. C.
5.2 Baselines and Evaluation Metrics
We compare DSPT5 with the following LLM baselines: Pre-trained LLaMA2 Touvron et al. (2023)  [^4] We implement the 7B version of LLaMA2., ChatGPT Brown et al. (2020)  [^5] https://chatgpt.com/, and Claude2 Anthropic (2023); Fine-tuned BART Lewis et al. (2020), T5 Raffel et al. (2020), and FLAN-T5 Chung et al. (2022). Crowd sampling is only added to DSPT5 for evaluation, as it is one component of DSPT5. All other LLMs baselines use their default decoding strategies. We evaluate DSPT5 with other SOTA LLMs using the following automatic evaluation metrics. (1) Embedding-based metrics: BERTScore (F1) Zhang* et al. (2020), document-level BLONDE (F1) Jiang et al. (2022), sentence-level and document-level BLEU Lin and Och (2004); (2) Word-based metrics: ROUGE1, ROUGE2 Lin (2004) and METEOR Banerjee and Lavie (2005); (3) End-to-end metrics: COMET Rei et al. (2022); (4) Simplicity: SARI Xu et al. (2016a).(5) Diversity: LCTR (Lexical Translation Consistency Ratio) Lyu et al. (2021).(6) Readability: FRES (Flesch Reading-Ease Score) Flesch (1979); (7) Toxicity Team et al. (2022) for safety evaluation.
5.3 Results and Analysis
5.3.1 Main Results
Table 2: Performance evaluation of all baselines with twelve different metrics. The abbreviations stands for Agriculture and Life Science, Architecture, Arts and Design, Engineering, Liberal Arts and Human Sciences, Natural Resources and Environment, Science, Business, and Veterinary Medicine.
Model ALS AAD ENG LAHS NRE SCI BUS VM ALS AAD ENG LAHS NRE SCI BUS VM
s-BLEU (%) d-BLEU (%)
BART 9.25 9.59 10.67 18.01 12.66 7.88 14.34 9.51 24.67 32.68 23.88 33.18 26.20 19.51 27.04 20.49
T5 7.93 11.48 8.39 16.63 11.91 7.33 15.35 8.33 25.61 32.79 24.82 33.58 23.80 17.00 23.77 19.56
FLAN-T5 8.89 12.27 8.46 17.09 6.94 7.84 14.94 8.33 18.39 28.35 17.45 31.31 17.69 15.88 30.65 18.66
Claude2 1.53 2.08 0.29 2.32 0.83 0.50 2.79 1.19 2.10 8.91 1.06 6.30 2..87 1.33 9.52 3.02
ChatGPT 4.53 5.74 4.34 6.25 4.08 4.20 8.17 4.51 15.39 15.59 13.75 16.99 14.11 12.84 17.20 13.62
LLaMA2 3.39 3.22 3.56 2.39 2.77 2.93 4.02 2.39 16.70 18.51 15.02 18.64 15.51 13.21 15.66 15.54
Ours 10.84 14.09 11.25 22.12 9.19 8.13 20.30 11.38 24.95 33.53 24.98 38.80 28.11 21.31 35.31 23.42
BERTScore (F1%) BLONDE (F1 %)
BART 85.12 82.83 84.05 86.09 85.58 82.74 86.59 83.94 15.04 45.15 9.07 44.85 15.22 8.02 38.91 13.64
T5 85.26 83.18 84.26 85.65 84.48 81.70 85.54 83.77 14.64 38.72 8.49 43.14 33.10 7.49 34.47 25.39
FLAN-T5 83.02 81.84 82.26 84.93 82.94 81.22 86.70 82.86 13.42 37.21 7.48 43.02 27.31 7.04 39.79 27.98
Claude2 79.48 80.02 77.05 80.95 79.96 77.43 83.25 79.01 10.41 21.13 2.75 17.39 10.37 3.06 21.62 4.94
ChatGPT 84.67 82.48 83.52 84.75 84.65 83.05 85.17 83.89 27.73 30.99 11.80 30.67 26.70 11.13 11.54 21.79
LLaMA2 83.58 80.97 82.46 83.66 83.31 81.46 83.14 82.74 11.38 37.39 8.25 32.77 30.56 7.63 30.01 10.82
Ours 85.48 83.70 84.41 87.27 85.51 82.90 87.97 84.20 36.75 44.11 9.62 48.51 36.60 8.22 35.53 30.69
ROUGE1 (%) ROUGE2 (%)
BART 55.25 49.86 51.26 57.92 57.55 47.49 57.78 50.33 27.83 27.06 25.98 36.55 31.16 20.42 33.75 22.38
T5 53.94 49.74 50.88 55.34 52.07 43.40 52.22 46.66 29.72 27.34 25.48 35.08 27.19 18.45 28.64 21.39
FLAN-T5 47.44 45.27 45.22 54.07 47.37 42.48 57.71 46.77 21.54 22.17 19.20 32.30 18.93 16.87 35.75 19.63
Claude2 28.47 36.13 21.95 36.44 33.23 23.13 44.04 29.75 5.42 9.43 2.53 9.47 6.34 3.14 14.63 5.03
ChatGPT 51.50 45.61 47.97 50.32 51.38 46.06 50.35 48.31 20.08 16.75 17.88 21.89 19.67 16.55 21.76 17.54
LLaMA2 44.43 38.30 40.24 39.92 43.22 38.20 41.52 42.49 24.82 23.30 21.86 27.59 24.92 19.37 23.01 22.47
Ours 54.48 50.84 51.73 59.67 56.47 47.14 60.75 50.28 28.58 27.55 27.02 40.24 30.83 21.97 39.61 24.63
METEOR (%) COMET (%)
BART 41.27 39.57 39.28 48.65 45.71 35.25 44.66 38.28 81.52 75.52 80.76 80.28 82.14 78.56 83.14 80.35
T5 40.34 40.84 39.84 46.86 40.09 31.77 40.04 33.81 80.45 74.77 80.37 76.19 75.66 70.69 77.69 73.18
FLAN-T5 32.97 35.62 33.52 46.04 34.64 31.36 47.86 33.85 77.14 73.26 78.03 78.50 80.02 75.65 82.52 79.04
Claude2 17.50 28.53 14.96 25.81 20.88 15.34 30.85 20.59 74.31 74.38 67.74 75.53 78.02 68.72 82.70 75.05
ChatGPT 38.94 36.86 37.47 41.68 39.71 36.06 40.56 36.09 85.49 80.30 84.00 83.29 85.79 82.98 85.29 84.10
LLaMA2 42.53 39.88 41.13 43.68 44.10 39.51 41.25 41.47 82.11 77.85 82.22 81.17 82.65 79.90 82.88 80.84
Ours 40.50 42.51 40.74 52.67 45.05 36.65 50.56 38.01 81.37 75.90 80.64 81.76 82.09 77.15 84.07 78.36
SARI LTCR (%)
BART 37.22 34.08 36.87 36.34 37.36 36.91 42.43 36.45 59.36 60.85 58.29 55.53 59.97 63.83 61.98 64.02
T5 37.89 35.89 37.88 36.83 36.11 37.21 47.75 36.92 57.61 60.82 56.18 52.27 60.51 62.74 58.84 66.01
FLAN-T5 36.73 34.12 36.78 35.43 34.83 37.76 48.62 36.71 61.41 63.91 60.63 56.89 64.59 84.12 61.19 67.16
Claude2 29.34 32.44 28.85 28.48 30.17 31.19 34.79 32.08 71.66 72.94 68.61 68.46 72.21 73.19 70.99 77.28
ChatGPT 37.83 35.41 36.90 34.50 36.61 38.53 40.07 38.07 58.87 65.38 57.61 59.64 61.14 61.97 64.28 63.35
LLaMA2 46.43 40.51 43.76 47.45 46.31 43.53 44.99 44.55 57.90 61.00 57.90 55.70 61.10 62.56 64.37 62.78
Ours 37.31 36.01 38.21 38.95 37.23 37.11 48.67 36.50 56.81 61.33 55.25 51.25 58.35 62.18 56.46 63.46
FRES Mean Toxicity (%)
BART 32.22 40.18 30.60 27.66 39.47 30.80 19.91 29.79 0.06 0.83 0.07 0.71 0.04 0.08 0.09 0.06
T5 32.83 41.60 31.01 27.96 40.69 31.62 20.72 31.82 0.12 0.74 0.08 1.00 0.04 0.11 0.10 0.06
FLAN-T5 40.89 40.89 31.41 29.48 40.99 32.63 22.34 31.62 0.07 0.54 0.08 1.40 0.04 0.11 0.08 0.05
Claude2 31.99 41.97 13.65 21.60 40.45 12.63 26.40 29.45 0.06 0.49 0.10 0.93 0.09 0.11 0.10 0.23
ChatGPT 42.11 42.41 33.85 31.72 41.70 34.15 29.69 42.21 0.03 0.06 0.05 0.32 0.03 0.05 0.09 0.03
LLaMA2 34.26 41.60 31.92 30.70 33.54 32.73 22.85 32.53 0.11 0.55 0.16 1.06 0.08 0.13 0.13 0.11
Ours 33.65 41.50 31.62 28.17 41.19 32.33 21.74 33.34 0.12 1.19 0.04 0.97 0.02 0.06 0.04 0.03
Table 3: Ablation study over DSPT5 components on VTechAGP across eight colleges. DSPT5dynamic is the model only using dynamic hard prompt from keyBERT. DSPT5dyn+soft is the model using dynamic soft prompt from the prompt encoder. DSPT5dyn+soft+con is the model fine-tuned using dynamic soft prompt with contrastive learning loss. DSPT5all is the final version of the model using dynamic soft prompt with contrastive learning for training and crowd sampling for inference.
Model s-BLEU d-BLEU BERTScore BLONDE ROUGE1 ROUGE2 METEOR COMET SARI LTCR FRES Toxicity
T5pre-trained 3.76 0.84 78.61 11.34 23.74 12.06 11.74 59.66 29.81 78.01 36.36 0.23
T5fine-tuned 6.02 19.14 82.14 22.14 45.53 22.14 36.44 68.74 36.55 62.77 31.54 0.20
DSPT5dynamic 8.78 22.83 83.83 28.43 49.77 25.33 38.35 77.28 36.71 60.19 32.31 0.24
DSPT5dyn+soft 9.87 25.79 84.43 34.92 51.14 26.98 40.64 78.91 37.03 59.11 32.34 0.11
DSPT5dyn+soft+con 9.88 26.78 84.58 33.99 51.66 26.82 40.90 78.54 37.03 59.05 32.33 0.13
DSPT5all 9.93 26.99 84.52 35.75 51.45 27.59 41.40 79.06 37.05 58.77 32.39 0.17
The experiment results on VTechAGP are shown in Table 2. We observe that: First, for the accuracy measurement between the generated text and the ground truth (i.e., embedding-based metrics and word-based metrics), our proposed model achieves the best performance in most cases. In addition, fine-tuned models (i.e. Bart, T5, FLAN-T5) always show better translation performance (BLEU, BERTScore, and ROUGE) than pre-trained LLMs (i.e. Claude2, ChatGPTP, LLaMA2). We believe this is because when the reference text (ground truth) is used to update the model’s parameters during the fine-tuning process, the model adapts to the nuances and characteristics of our specific academic to general audience text paraphrasing task. In contrast, the pre-trained LLMs are trained on large and diverse datasets and are not optimized for the academic domain. This indicates that it is possible to fine-tune a lightweight language model that can achieve competitive translation performance with LLMs on the high-quality VTechAGP dataset.
Second, pre-trained LLMs outperform fine-tuned lightweight models in end-to-end metrics, simplicity, diversity, and readability. However, different LLMs show distinctive strengths and capabilities. For example, LLaMA2 performs well at text simplification. Claude2 is only good at generating diverse text, and ChatGPT demonstrates the best performance for end-to-end metrics and readability. Each LLM has a unique architecture and has been pre-trained on different datasets, which may explain the observed differences. Finding an LLM that performs well in all evaluation metrics across all domains is challenging, indicating that there is still much room for improvement in LLMs for academic-to-general-audience text paraphrasing tasks. Besides, we need to further explore new evaluation metrics that can more accurately reflect the quality of the generated general audience text.
5.3.2 Ablation Study
To assess the performance of each component in DSPT5, we conducted the ablation study. For each run, we randomly sampled 15 data points from each college to construct a balanced dataset for the ablation study (see ablation study setup in Appendix C). The results are presented in Table 3 as the mean value of five different subsets. We observe that: (1) Fine-tuning the LLM can significantly improve its performance. This is because fine-tuning the model on a customized dataset can enable it to learn domain-specific nuances and identify task-specific patterns. (2) Adding the dynamic soft prompt consisting of domain-specific keywords can further improve the performance. The results verify that adding the extracted academic keywords with the phrase "reduce the domain knowledge of" in the prompt template while training with contrastive-generative loss can reduce the domain knowledge. Besides, the crowd sampling decoding is only implemented in the inference stage, it is worth adding the module for further performance improvement; (3) The diversity LTCR and readability FRES decrease after fine-tuning T5. We suspect that this is because T5 is pre-trained on diverse and large datasets, resulting in the generated text being represented by a diverse vocabulary. Moreover, the output of T5pre-trained is always shortened compared to T5fine-tuned, making the output easier to understand (higher FRES). Thus, new evaluation metrics for academic-to-general-audience text paraphrasing tasks are needed in the future.
5.3.3 Human Evaluation
Table 4: Mean human evaluation ratings 1-5 (the higher the better) of different models on VTechAGP. Reported from left to right are: comprehensiveness, layness, meaning preservation, conciseness and fluency.
COM LAY MP CON FLU
BART 3.83 2.83 3.67 3.58 3.83
T5 3.61 2.95 3.35 3.5 3.51
FlanT5 3.71 3.13 3.41 3.66 3.76
Claude2 3.08 2.95 2.72 3.55 3.26
ChatGPT 4.17 3.66 3.83 3.71 4.31
LLaMA2 4.05 3.61 3.73 3.80 4.31
Ours 4.17 3.57 3.90 3.87 3.95
The human evaluation setup is discussed in detail in Sec. F in the Appendix. Human evaluation results are shown in Table 4. From Table 4, we observe that generally ChatGPT, and our proposed DSPT5 show better performance than other baselines. Based on the observations, new information with simpler words is always introduced in the generated outputs from ChatGPT. Those sentences with simpler words make the generated abstracts easier to understand, resulting in a better layness. However, it also significantly changes the meaning of the original sentences and makes the sentences longer with redundency. That’s why ChatGPT shows the best performance in layness, while not the best in meaning preservation and conciseness. DSPT5 attempts to paraphrase the abstracts with non-technical words while retaining the original meaning as closely as possible. While keeping the content concise, it may sacrifice fluency between sentences. Although ChatGPT shows competitive results, it is NOT an open-sourced model, which may raise concerns about data security and high cost issues. Alternatively, the DSPT5 can be fully controlled and owned with only a single V100 GPU for fine-tuning jobs for the academic-to-general-audience text paraphrasing task.
6 Conclusion
We created VTechAGP, an academic-to-general-audience document-level translation benchmark dataset, which consists of 4,938 academic and general audience abstract pairs with their corresponding auxiliary information such as title, subject terms, etc. We explore several SOTA LLMs to establish the baseline performance on the text paraphrasing task. We propose a new model, DSPT5, which includes dynamic soft prompt generation with hybrid loss during the training phase and a new crowd-sampling strategy in the inference stage. We evaluate the datasets and models from the perspective of document-level embedding-based, word-based, end-to-end metrics, simplicity, diversity, readability, toxicity and human evaluation. Extensive experimental results on 8 different colleges of VTechAGP show that DSPT5 achieves comparable results with other SOTA benchmark models.
7 Limitations
Based on the exploration of the datasets in Sec. B and the experiments of the benchmark models in Sec. 5, we identify the following limitations (remaining challenges) and future directions:
(1) Dataset: Although VTechAGP is a recent and distinctive dataset, it suffers in terms of size, with a total of 4,938 document pairs. This is due to the fact that the policy of including a general audience version of the abstract for all theses and dissertations submitted to VTechWorks began to be implemented in 2016. However, this aspect of the limitation should be gradually overcome as VTechWorks continues to collect and maintain this collection of documents. With the increase in graduate student enrollment at Virginia Tech in recent years, it is safe to assume that VTechAGP will continue to grow in the coming years. In addition, VTechAGP consists of pairs of abstracts carefully written by students, encapsulating their work during their time in graduate school. With about 50% of the corpus composed of Ph.D. dissertations and the other half composed of M.S. theses, we believe that VTechAGP should achieve good quality due to the amount of time and effort put into the creation process. As the size of the dataset grows, VTechAGP will become more robust for text paraphrasing tasks in the near future.
(2) Model: While transformer-based LLMs retain their popularity for text generation tasks, our experiment results in Sec. 5 show that there is still a huge room for further improvement of LLMs in academic to general audience text paraphrasing tasks. With the rapid development of LLMs, more robust and efficient LLMs are emerging. For example, the recently released model Gemini Team et al. (2023) shows great potential and competitive results compared to ChatGPT and LLaMA2 in natural language understanding-related tasks. In addition, some recently released lightweight decoder-only models (e.g., Phi-2 Li et al. (2023)) also exhibit great potential for text generation tasks. As the VTechAGP dataset grows, it will be able to support fine-tuning for more powerful open-source LLMs (i.e. Phi-2, Gemini).
(3) Evaluation Metric: Although we have implemented twelve different automatic evaluation metrics, including document-level embedding-based metrics, word-based metrics, end-to-end metrics, simplicity, diversity, readability, and toxicity, we still lack a representative metric that can evaluate the quality of the generated text in terms of general audience understanding while remaining scientifically accurate and easy to comprehend. The existing evaluation metrics for simplicity such as SARI Xu et al. (2016a) focus on measuring the goodness of words that are added, deleted, and kept by the system. In addition, regarding the formula of FRES Flesch (1979), such readability evaluation metric only considers the number of words and the number of sentences when evaluating the ease of understanding for children of different grades. Instead of just simplifying sentence structure and using words with fewer letters, our task aims to reduce the domain knowledge of the generated text while remaining scientifically correct, so that the general audience or people/researchers in other research fields (departments) can still understand the core idea for further interdisciplinary collaboration. New automatic evaluation metrics or large-scale human evaluation approaches for academic-to-general-audience text paraphrasing tasks need to be explored in future work.
Appendix A Dataset Comparison
In this section, we explore different features (i.e. document-level or sentence-level, multiple domains or single domain, with auxiliary information or not, in English or other languages) of existing public datasets for text simplification and text paraphrasing, which motivated us to collect and publish a new academic-to-general-audience text paraphrasing dataset VTechAGP, which is a document-level, multi-domain English text paraphrasing dataset with domain-specific auxiliary information. We show the comparison of our proposed dataset VTechAGP with other existing public datasets. The details of features are displayed in Table 5.
Table 5: Comparison between VTechAGP and other text simplification or paraphrase datasets. Columns indicate whether the dataset has the feature.
Doc. Multi Domains Side Info. Eng.
Newsela Xu et al. (2015)
WikiSmall Zhu et al. (2010)
WikiLarge Zhang and Lapata (2017)
WikiAuto Jiang et al. (2020)
TurkCorpus Xu et al. (2016b)
ASSET Alva-Manchego et al. (2020)
MedLane Luo et al. (2022b)
OneStopEnglish Vajjala and Lučić (2018b)
Lexica Hewett and Stede (6)
DEplain Stodden et al. (2023)
Quora Wang et al. (2017)
Paralex Fader et al. (2013)
MRPC Dolan and Brockett (2005)
STS Cer et al. (2017)
SNLI Bowman et al. (2015)
ParaNMT-50M Wieting and Gimpel (2017)
MSCOCO Lin et al. (2014)
VTechAGP (ours)
Table 6: Description of CSV File Columns in the Dataset
Column Name Description
identifier_uri Persistent identifier (CNRI handle)
title Title
abstract Regular abstract
abstract_general General audience abstract
subject_terms List of subject terms
discipline Field of study for the degree awarded
department Name of the academic department
degree Degree awarded
degree_level Degree level (‘doctoral’ or ‘masters’)
type Type of ETD (‘thesis’ or ‘dissertation’)
Table 7: An example of academic abstract and general-audience abstract pair.
Academic Abstract
Breath-first search (BFS) is a fundamental building block in many graph-based applications, but it is difficult to optimize for a field-programmable gate array (FPGA) due to its irregular memory-access patterns. Prior work, based on hardware description languages (HDLs) and high-level synthesis (HLS), address the memory-access bottleneck of BFS by using techniques such as data alignment and compute-unit replication on FPGAs. The efficacy of such optimizations depends on factors such as the sparsity of target graph datasets. Optimizations intended for sparse graphs may not work as effectively for dense graphs on an FPGA and vice versa. This thesis presents two sets of FPGA optimization strategies for BFS, one for near-hypersparse graphs and the other designed for sparse to moderately dense graphs. For near-hypersparse graphs, a queue-based kernel with maximal use of local memory on FPGA is implemented. For denser graphs, an array-based kernel with compute-unit replication is implemented. Across a diverse collection of graphs, our OpenCL optimization strategies for near-hypersparse graphs delivers a 5.7x to 22.3x speedup over a state-of-the-art OpenCL implementation, when evaluated on an Intel Stratix 10 FPGA. The optimization strategies for sparse to moderately dense graphs deliver 1.1x to 2.3x speedup over a state-of-the-art OpenCL implementation on the same FPGA. Finally, this work uses graph metrics such as average degree and Gini coefficient to observe the impact of graph properties on the performance of the proposed optimization strategies.
General-Audience Abstract
A graph is a data structure that typically consists of two sets – a set of vertices and a set of edges representing connections between the vertices. Graphs are used in a broad set of application domains such as the testing and verification of digital circuits, data mining of social networks, and analysis of road networks. In such application areas, breadth-first search (BFS) is a fundamental building block. BFS is used to identify the minimum number of edges needed to be traversed from a source vertex to one or many destination vertices. In recent years, several attempts have been made to optimize the performance of BFS on reconfigurable architectures such as field-programmable gate arrays (FPGAs). However, the optimization strategies for BFS are not necessarily applicable to all types of graphs. Moreover, the efficacy of such optimizations oftentimes depends on the sparsity of input graphs. To that end, this work presents optimization strategies for graphs with varying levels of sparsity. Furthermore, this work shows that by tailoring the BFS design based on the sparsity of the input graph, significant performance improvements are obtained over the state-of-the-art BFS implementations on an FPGA.
Appendix B Dataset Analysis
VTechAGP contains a total of 4,938 document pairs (academic abstract and general audience abstract), where 52.8% of the documents are from Ph.D. dissertations and 47.2% of the documents are from Master’s theses. Table 1 shows the data statistics of VTechAGP across eight different colleges. VTechAGP is divided into eight colleges because data from different colleges exhibits distinct statistics. In addition, the dataset divided into colleges can also support potential cross-domain research tasks and provide the convenience for researchers focusing on a specific domain in the future. The document pairs are categorized based on colleges since authors from the same college may share some common terminology and background knowledge. Therefore, these categories should exhibit similar statistics. Table 1 reports some basic dataset statistics, including the number of documents for each college, the average number of sentences for each document, and the average sentence length for the source and target texts, respectively. We also report textual and lexical diversity (MTLD), which reflects the average number of words in a row for which a certain type-token ratio is maintained McCarthy (2005). To better illustrate the difference in readability between the source and target texts, we report Readability Consensus, [^6] https://pypi.org/project/textstat/ which is a combined evaluation of Flesch Kincaid Grade, Flesch Reading Ease, SMOG Index, Coleman-Liau Index, Automated Readability Index, Dale-Chall Readability Score, Linsear Write Formula, and Gunning FOG Formula. From Table 1, we observe that the general audience abstracts have fewer sentences and shorter sentence lengths compared to the academic abstracts. Furthermore, academic abstracts have better lexical diversity than general audience abstracts, except for the College of Architecture, Arts and Design, and the College of Business. In general, academic abstracts have a higher level of comprehension difficulty compared to general audience abstracts, except the College of Engineering. Analysis for the College of Engineering, which represents 49.6% of the total dataset, is discussed below.
Table 8: Dataset statistics of VTechAGP over 13 departments in the College of Engineering. Statistics are reported in the format of academic (source) / general audience (target) doc.
Department # Doc. Readability Consensus
Aerospace Eng. 181 14th-15th / 14th-15th
Biomedical Eng. 234 13th-14th / 10th-11th
Building Construction 19 15th-16th / 15th-16th
Chemical Eng. 46 16th-17th / 11th-12th
Civil Eng. 234 16th-17th / 15th-16th
Computer Sci. 392 14th-15th / 15th-16th
Electrical & Computer Eng. 749 15th-16th / 15th-16th
Eng. Education 31 17th-18th / 16th-17th
Environmental Sci. & Eng. 23 15th-16th / 14th-15th
Industrial & Systems Eng. 86 16th-17th / 15th-16th
Materials Sci. & Eng. 72 16th-17th / 11th-12th
Mechanical Eng. 349 15th-16th / 14th-15th
Mining Eng. 35 15th-16th / 15th-16th
The distinct category (College of Engineering) contains the largest number of document pairs, representing 49.6% of the total dataset, due to the large number of students from 13 different departments in the College of Engineering. To better understand the characteristics of this category, we divided the documents in the College of Engineering based on their corresponding departments. Details are given in Table 8. Thus, we expect document pairs within this college to be more diverse compared to others. From Table 8, we observe that documents from most departments in the College of Engineering have a similar readability consensus between academic abstracts and general audience abstracts. The only outlier is the Department of Computer Science, which shows that general audience abstracts are more difficult to understand than academic abstracts. On the other hand, the Department of Electrical and Computer Engineering alone represents 15.2% of the total dataset, while maintaining an above-average readability consensus for its general audience abstracts, thus contributing to the higher readability consensus for the College of Engineering. We conjecture as follows: 1) Various complex terminologies from some research fields are inevitably retained in general audience abstracts because they cannot be replaced by simple terms. 2) Concepts in certain areas of engineering research are difficult to explain or rewrite into short and concise sentences. 3) Some engineering students are not as good at simplification writing as students from other colleges.
Appendix C Parameter Settings
For the hyperparameters and configuration of DSPT5, we implement DSPT5 in PyTorch and optimize it with the AdamW optimizer. For pre-trained LLMs, we retrieve the generated output through their API by giving the prompt "Generate another version of the provided document for the general audience". We use grid search to tune the hyperparameters. The learning rate is $\in\left\{5e-2,5e-3,5e-4,5e-5\right\}$. The contrastive loss weight $\lambda$ in Equ. 8 and the alignment function weight $\gamma$ in Equ. 10 are chosen from 0.1 to 0.9 with a step size of 0.2. The number of candidates $C_{n}$ in Sec. 4.3 is chosen from $\left\{4,8,16\right\}$. The batch size is 4. The length of the source text, fixed prompt template, and keywords are 512, 16, and 16, respectively. The fixed prompt template used for baselines is: ‘Generate another version of the provided document for general audiences.’. We choose the hyperparameters based on the validation set. Because the dataset is not a balanced dataset, where College of Engineering takes up most of the datasets. We split a validation set in the domain of College of Engineering documents. This is because as shown in Table 1, it is not realistic to have a validation set in the Business domain, which has only 63 documents in total. The size of the dataset will only grow as new master’s theses or PhD dissertations are submitted. In the future, when the dataset is large enough, we will split out a validation set that includes data from all colleges. The final learning rate we used is 5e-5. However, this can be changed according to different batch sizes when fine-tuning the model. The parameters we used are searched by a validation set from the engineering college documents (which takes the highest percentage of the total documents). The contrastive loss weight is 0.3, the alignment function weight is 0.1, and the number of candidates is 16. If people want to work in a different domain (other than College of Engineering), different parameters may have different performances. For the temperature in contrastive loss, we use the InfoNCE loss from van den Oord et al. (2018), and we use the default temperature value of 0.1. The temperature in decoding is 0.5. For model details shown in Figure 1, we use the same T5-encoder that shares parameters to keep our model DSPT5 as small (efficient) as possible.
For ablation study setup, the model is trained on the entire training dataset without sampling. We only do the sampling on the evaluation dataset for the ablation study. Since the data split is 0.8:0.2 on each college domain, the evaluation set is also not a balanced dataset. In total, there are about 988 documents in the evaluation set, which is used for the evaluation in Table 2. After sampling, we have around 120 documents (which is now a balanced evaluation set covering all colleges) for the evaluation set ONLY used for the ablation study Table 3. We reported the average performance including all colleges in Table 3 for ablation study. Since the College of Engineering takes up almost half of the data shown in Table 1, the result of the average performance will be heavily influenced by the performance of the College of Engineering. Therefore, we sampled a balanced evaluation set to report the ablation study.
Appendix D More Results
In this section, we provide the combined test set results of all eight colleges, which is shown in Tabel 9. From Table 9, we observe that our proposed DSPT5 outperforms all other baselines in the evaluation metric of s-BLEU, d-BLEU, BERTScore, BLONDE, ROUGE and METEOR. The closed-source model ChatGPT shows better performance in system-level evaluation metrics such as COMET. We conjecture that this is probably because such pre-trained LLMs (i.e. ChatGPT, Claude2, LLaMA2) are pre-trained on large quantities of documents, which make these models generate text more fluently, resulting in better system-level evaluation metrics. Alternatively, our proposed model DSPT5 is fine-tuned on our customized dataset VTechAGP, which results in better performance in word-level evaluation metrics.
Table 9: Combined test results of all eight colleges.
s-BLEU d-CLEU BERTScore BLONDE ROUGE1 ROUGE2 METEOR COMET SARI LTCR FRES Toxicity
BART 11.24 25.09 84.28 16.82 52.33 27.15 40.58 80.27 36.79 59.19 31.12 0.19
T5 9.75 25.16 84.15 17.20 50.63 26.30 39.62 77.71 37.56 57.44 31.73 0.24
FLAN-T5 9.77 19.93 82.64 16.26 46.57 21.23 35.24 77.66 36.65 63.99 32.98 0.28
Claude2 0.88 2.50 78.28 7.03 26.44 4.57 18.04 70.78 29.62 70.16 19.57 0.23
ChatGPT 4.70 14.36 83.77 17.65 48.45 18.52 38.07 83.84 36.86 59.40 35.33 0.08
LLaMA2 3.21 15.71 82.58 14.70 40.43 22.83 41.49 81.54 44.51 58.78 32.50 0.29
Ours 12.41 27.02 84.75 21.11 52.71 28.59 42.15 80.21 38.01 56.50 32.32 0.23
Appendix E Case Study
Table 10 shows an example of general audience abstracts generated by different LLMs on VTechAGP (Department of Chemical Engineering). The first few sentences of the document are displayed. Fine-tuned T5 and FLAN-T5 show similar outputs with source text, so we have not included them in Table 10. The key idea of the source text is: Polymers are important in life and its manufacturing process is critical to industry. We observe that the pre-trained LLMs can improve the readability to some extent. For example, ChatGPT and LLaMA2 rewrite ‘automotive industry’ as ‘cars’. Claude2 translates ‘Ziegler-Natta catalysts’ to ‘a specific type of catalyst’. However, some academic domain-specific words (technical terminology) still exist. For example, the generated outputs still contain the phrases ‘high-density polyethylene’ and ‘linear low-density polyethylene’ from Claude2 and ChatGPT, ‘Ziegler-Natta catalysts’ from ChatGPT and LLaMA2. Despite the catalyst terminology, LLaMA2 summarizes the academic text best, which is consistent with the experimental results of the simplicity metric (SARI) in Table 2. In contrast, the fine-tuned generative language models (i.e., BART and Ours) do not introduce these technical terminologies. Fine-tuned BART generates some unrelated phrases (i.e. kinetic parameters), whereas our proposed model DSPT5 can still express the key idea from the source text that polymers are important and widely used in life and their manufacturing process is critical to industry. At the same time, there is no technical terminology in the output text generated by DSPT5. We did not include the output of fine-tuned T5 and FLAN-T5 as they yield similar outputs to the source text. We speculate that this is because there is no appropriate domain-specific prompt provided for fine-tuning T5 and FLAN-T5 with the limited training data.
Appendix F Human Evaluation
To provide a comprehensive assessment of the generated general-audience abstract, we conducted a human evaluation involving our proposed model DSPT5 and all other baseline models using three independent experts  [^7] All judges have experience in scientific research and hold bachelor, master, and doctorate degree respectively.. Specifically, following a similar setting as Liu et al. (2024); Li et al. (2024); Song et al. (2024); Kew et al. (2023); Devaraj et al. (2021), our evaluation uses a random sample of 20 abstracts (10 in Computer Science and 10 in all other fields) from the test split VTechAGP considering the workload. Judges are presented with both the academic abstract and generated general-audience abstracts from seven models (DSPT5 and six baselines) for each data sample in total of 140 abstracts. Using a 1-5 Likert scale, the judges are asked to rate the model output based on five criteria: comprehensiveness, layness, meaning preservation, conciseness, and fluency. Details are shown in guidelines, which clarifies the meanings for each criteria in Figure 3 in the Appendix. Table 4 presents the average ratings from our human evaluation.
Appendix G Methodology Explanation Details
In Sec. 4, we discuss the contrastive loss infoNCE derived from different learning representations. Here we provide a detailed explanation. To textual level, $r$ represents a set of keywords (here a set of keywords are textual keywords, and then $r$ is the representation (embedding) of the several textual keywords). As we said, $r^{pos}$ and $r^{neg}$ are not sets of vectors. Instead, $r^{pos}$ is a vector and $r^{neg}$ is a vector. They are the vector representation of a set of textual keywords. If there are $m$ keywords in the golden label and keyBERT returns $n$ keywords (where $m<n$), we sort the keywords by confidence and select the top $m$ keywords. Then we concatenate the words for dot production. For example if there are two terms in the label, we select two positive keywords from KeyBERT. The selection is based on the ranking of the confidence value provided by KeyBERT. Although dot production is sensitive to the order of the keywords, currently we do a simple concatenation and we leave the works for exploring new ranking approaches to make the keyword extraction module extract closer or more relevant words to the golden label in the future as the main contributions for this paper is about the new dataset VTechAGP and implementations for benchmark models.
Appendix H Ethical Statements
The VTechAGP dataset, derived from theses and dissertations available in VTechWorks, [^8] Virginia Tech Electronic Theses and Dissertations: https://hdl.handle.net/10919/5534 is shared with a clear understanding of the ethical implications. Virginia Tech, as the institution where these ETDs were submitted, holds the right to distribute the content within the bounds of academic sharing and research purposes. The university’s policy ensures that the distribution of such academic materials aligns with educational objectives and respects the authors’ intellectual contributions.
Descriptive metadata, including titles, degree types, and departmental affiliations, is generally not subject to copyright restrictions. This metadata is typically viewed as factual information, which falls outside the scope of copyright protection. This understanding allows for the broad dissemination and analysis of metadata in academic and research contexts without infringing upon copyright laws. The abstracts, however, may present a different scenario. While the university facilitates their distribution through VTechWorks, the copyright of the abstract text itself could potentially remain with the authors. Nevertheless, the use of abstracts in the VTechAGP can be justified under the principles of fair use, particularly given the dataset’s purpose in advancing scholarly research, education, and understanding. The dataset does not replicate the full ETDs but instead provides a means to study and analyze the structure and nature of academic abstracts themselves, thereby contributing to the field of information retrieval and digital libraries.
Table 10: Examples of academic to general audience text paraphrasing comparison. The target text shows the gold label.
Source Text Polyolefins are one of the most widely used commodity polymers with applications in films, packaging and automotive industry. The modeling of polymerization processes producing polyolefins, including high-density polyethylene (HDPE), polypropylene (PP), and linear low-density polyethylene (LLDPE) using Ziegler-Natta catalysts with multiple active sites, is a complex and challenging task…
Target Text Almost everything we see around us from furniture, electronics to bottles, cars, etc. are made fully or partially from plastic polymers. The two most popular polymers which comprise almost two-thirds of polymer production globally are polyethylene (PE) and polypropylene (PP), collectively known as polyolefins. Hence, the optimization of polyolefin manufacturing processes with the aid of simulation models is critical and profitable for chemical industry.
BART Polyolefins are one of the most widely used commodity polymers with applications in films, packaging and automotive industry. The modeling of polymerization processes producing polyfins involves many processes and their kinetic parameters that have been validated by scientific institutions and market models…
Claude2 Polyolefins are a category of plastics that are utilized across many industries for products such as packaging materials, plastic bags, and automobile parts. Modeling the manufacturing processes that create popular polyolefins like high-density polyethylene, polypropylene, and linear low-density polyethylene, when using a specific type of catalyst, involves intricate and difficult steps…
ChatGPT Polyolefins are commonly used plastics found in items like films, packaging materials, and parts for cars. Creating these plastics, such as high-density polyethylene (HDPE), polypropylene (PP), and linear low-density polyethylene (LLDPE), through polymerization processes involving Ziegler-Natta catalysts with multiple active sites is a complex and challenging undertaking.
LLaMA2 Polyolefins are common plastics used in packaging, films, and car parts. Modeling how these materials are made is complex and challenging, especially when using Ziegler-Natta catalysts with multiple active sites…
Ours The global population continues to grow, and with it, the demand for polyolefin will further increase. Thus, there will never be a non-polymerization process, and thus, many parts of the global production chain must be produced synthetically. Polyolefin is the most widely used commodity in the automotive and aerospace industries. The manufacturing of polyolefins is challenging since…
Figure 3: Human evaluation guideline.